OpenStudy (wintersuntime):

Solve for this equation 3(x-2)=9

OpenStudy (photon336):

do you know how to use the distributive property?

OpenStudy (wintersuntime):

you multiply 3 to x and 2 right ?

OpenStudy (photon336):

\[3(x-2) = 3*x-?\]

OpenStudy (photon336):

Can you see the pattern?

OpenStudy (wintersuntime):

yes so it would be 3x-3*2=9

OpenStudy (wintersuntime):

then 3x-6=9 ??

OpenStudy (wintersuntime):

hello ?

OpenStudy (wintersuntime):

@Photon336

OpenStudy (photon336):

:) yes

OpenStudy (wintersuntime):

Is that right ?

OpenStudy (photon336):

Yep you're right. now, to solve for x we need to get x by itself

OpenStudy (wintersuntime):

How do I do that I forgot xD

OpenStudy (photon336):

okay so we have 3x-6 = 9 right? first step would be to add 9 to both sides 3x-6+6 = 9+6 3x = 15

OpenStudy (wintersuntime):

Then divide ?

OpenStudy (wintersuntime):

3x

OpenStudy (wintersuntime):

x=5